# Intelligent Polarforming Antenna Enhanced Sensing and Communication: Modeling and Optimization

### v1

@@ Abstract
In this paper, we propose a novel intelligent polarforming antenna (IPA) to achieve cost-effective wireless sensing and communication. Specifically, the IPA can enable polarforming by adaptively controlling the antenna’s polarization electrically as well as its position/rotation mechanically, so as to effectively exploit polarization and spatial diversity to reconfigure wireless channels for improving sensing and communication performance. To analyze the performance gain of IPA, we study an IPA-enhanced integrated sensing and communication (ISAC) system that utilizes user location sensing to facilitate communication between an IPA-equipped base station (BS) and IPA-equipped users, by focusing on a new practical channel setup where the locations of users are nearly time-invariant but their orientations may change frequently (e.g., mobile phones rotated by spectators seated in a stadium while taking live photos). First, we model the IPA channel in terms of transceiver antenna polarforming vectors and antenna positions/rotations. We then propose a two-timescale ISAC protocol, where in the slow timescale, user localization is first performed, followed by the optimization of the BS antennas’ positions and rotations based on the sensed user locations; subsequently, in the fast timescale, transceiver polarforming is adapted to cater to the instantaneous orientation of user devices in three-dimensional (3D) space, with the optimized BS antennas’ positions and rotations. We propose a new polarforming-based user localization method that uses a structured time-domain pattern of pilot-polarforming vectors to extract the common stable components in the IPA channel across different polarizations based on the parallel factor (PARAFAC) tensor model. Moreover, we maximize the achievable average sum-rate of users by jointly optimizing the fast-timescale transceiver polarforming, including phase shifts and amplitude variations, along with the slow-timescale antenna rotations and positions at the BS. Simulation results validate the effectiveness of polarforming-based localization algorithm and demonstrate the performance advantages of polarforming, antenna placement, and their joint design in comparison with various benchmarks without polarforming or antenna position/rotation adaptation.
@@
Intelligent polarforming antenna (IPA), polarforming optimization, antenna position/rotation optimization, integrated sensing and communication (ISAC), location sensing.
@@ I Introduction
To enable the transceiver’s full flexibility in antenna polarization and placement, we propose a novel intelligent polarforming antenna (IPA), as a new approach to improve the performance of wireless sensing and communication without the need to add more antennas and bear their additional cost and energy consumption. Specifically, an IPA tunes the polarforming vector with adjustable signal amplitude and phase to control the polarization of each antenna at the transmitter and/or receiver, thus allowing dynamic adjustment of antenna polarization in response to real-time channel conditions and depolarization effects by exploiting polarization diversity. This is achieved through the electronically tunable polarformer, which contains phase shifters and attenuators that collectively adjust the phase and amplitude of the transmitted/received signal (see Fig. 1). Furthermore, each IPA can be independently adjusted in terms of position and/or rotation to adapt to the channel spatial distribution. This can be realized through mechanical control, which utilizes external mechanical structures with the aid of actuators, such as electric motors and precision gears, to achieve physical movement (see Fig. 1).
Figure 1: Illustration of the IPA-enhanced sensing and communication system.
@@
It is worth noting that the IPA system proposed in this paper differs significantly from our recently proposed six-dimensional movable antenna (6DMA) [9, 10, 11] and traditional polarization reconfigurable antenna [12]. Firstly, a mechanically controlled 6DMA, consisting of multiple three-dimensional (3D) rotatable and positionable antennas/subarrays, either lacks the antenna polarization control or is designed under a fixed polarization condition [13, 14, 15, 16, 17]. Secondly, although traditional polarization-reconfigurable antennas can adjust antenna polarization, they tune only the amplitude of the signal, which does not fully exploit polarization diversity. Moreover, polarization reconfigurable antennas with fixed antenna position/rotation cannot leverage antenna movement to achieve additional spatial DoFs. In contrast, the IPA with polarformer aims at collectively adjusting both the amplitude and phase of the signals to fully explore the benefits of polarization diversity, as well as adjusting its antenna position/rotation to enhance spatial DoFs. A summary of the above comparison is given in Table I.
@@
Besides enhancing communication performance, the proposed IPA is to help improve wireless sensing/localization performance. Various wireless localization systems exist, such as those based on WLAN or cellular networks, which can determine the location of a user or target object using dedicated wireless reference signals between reference nodes and target nodes [1, 18, 19, 20]. A target object at an unknown location is often called agent node, while a reference node with a known position is referred to as anchor node. In some cases, the agent node determines its location by receiving reference signals from anchor nodes. However, there are other cases where the anchors receive reference signals from agents to localize the agents, and then inform them the estimated locations via a separate communication link [21]. Here, we consider the latter case, where multiple IPAs are deployed at the anchor/base station (BS) to receive reference signals from single-IPA users for user localization.
@@
Motivated by the above, we consider an IPA-enhanced ISAC system that utilizes user location sensing to facilitate communication between IPA-equipped users and an IPA-equipped BS. In our considered system, the transmit and receive polarforming vectors, along with the positions and rotations of the IPA subarrays at the BS, need to be jointly designed to achieve optimal sensing and communication performance. However, the design of these parameters depends on the user’s position and orientation. Thus, we consider a practical two-timescale channel setup, where the users’ positions change slowly and can be assumed to be fixed within a sufficiently long period termed as location coherence time, while their orientations may vary frequently during this period (intentionally or non-intentionally). This channel setup is applicable to many scenarios, such as spectators seated in a stadium watching a football match, where their smartphones remain stationary in position but are frequently rotated to take photos in different directions (see Fig. 2 (a)). Another example is for users playing virtual reality (VR) games at fixed locations, where their VR gadgets rotate frequently to enhance the gaming experience (see Fig. 2 (b)).
TABLE I: Comparison of IPA with Other Related Technologies.
@@
Intelligent polarforming antenna (IPA) High (Both phase and amplitude-adjustable polarforming) High (rotation and position) Electrical and mechanical Very high Medium
@@
In the proposed IPA-enhanced ISAC system, the positions and rotations of antennas are mechanically controlled, which may have limited movement speeds, rendering them more suitable for adapting to slow-timescale channel variations, e.g., user locations. In contrast, transmit and receive polarforming of IPA is electronically controlled, thus allowing their rapid adjustments to adapt to fast-timescale channel variations caused by, e.g., users’ frequent rotations. To design such a system, the acquisition of channel state information (CSI) is essential. However, since the BS can only measure user-BS channels at the positions and rotations where BS’s subarrays are physically deployed, it is difficult to obtain the instantaneous CSI across the entire antenna movement region at the BS and for arbitrary users’ positions and rotations. Fortunately, in our considered scenarios (see Fig. 2), the user’s locations remain unchanged within each location coherence time, which can be acquired and used for the optimization of BS antenna positions/rotations. Then, with the optimized antenna positions/rotations at the BS, the instantaneous CSI of users due to their rotations can be measured for designing the fast-timescale transmit and receive polarforming. The main contributions of this paper are summarized as follows.
@@
  • First, we propose the new IPA to electrically adjust antenna polarization through polarforming vectors to align the transmitter and receiver polarizations, as well as mechanically adjust antenna position/rotation to adapt to user spatial channel distribution. Moreover, we model the IPA channel and decouple the channel into stable unpolarformed components and dynamic polarformed components, which provides essential insights into our considered channel setup and facilitates the determination of parameters in the IPA-enhanced system.
@@
  • Next, we study an IPA-enhanced ISAC system that utilizes user location sensing to facilitate communications between IPA-equipped users and an IPA-equipped BS. For the new practical channel setup where user locations are nearly static but orientations may vary frequently, we propose a two-timescale ISAC protocol. In the slow timescale, user localization is first performed, based on which the BS antennas’ positions and rotations are optimized, while subsequently in the fast timescale, transceiver polarforming is adapted to cater to the instantaneous user/device rotations in 3D space.
@@
  • Moreover, we design a new polarforming-based user localization method that leverages controllable polarforming vectors to create a time-domain pattern of pilot-polarforming vectors. This enables the extraction of stable line-of-sight (LoS) components in the IPA channel across different polarizations using the parallel factor (PARAFAC) tensor. Thereby, user locations can be accurately determined from the estimated LoS components. Simulation results validate the effectiveness of the proposed localization method by exploiting the new DoF of polarforming.
@@
  • Finally, we formulate an optimization problem to maximize the weighted sum-rate of users by jointly designing transceiver polarforming and BS antenna positions/rotations based on sensed user locations. To solve this non-convex problem, we propose an efficient algorithm that accounts for IPA movement constraints, discrete polarforming amplitude and phase shifts, and enables efficient variable updates via closed-form solutions or simple iterations. We show by simulation that polarforming optimization improves the achievable rate over fixed polarization, with more gains achieved when being jointly optimized with antenna position/rotation.
@@
Figure 2: Examples of user rotation scenarios in IPA-enhanced ISAC systems.
@@
The rest of this paper is organized as follows. Section II presents the IPA architecture and the corresponding channel model. Section III presents a two-timescale protocol design for the IPA-enhanced ISAC system. Section IV presents the polarforming-based localization algorithm. Section V addresses antenna polarforming and position/rotation optimization. Section VI presents simulation results for performance evaluation. Finally, Section VII concludes this study.
@@ II System Model
We consider an IPA-enhanced ISAC system as shown in Fig. 1, where multiple IPAs are deployed at the BS to receive reference signals from single-IPA users for user localization as well as to communicate with these users over a given frequency band. Without loss of generality, we consider the downlink communications from the BS to users. In this section, we first describe the new IPA architecture and then present its corresponding channel model, which lays the foundation for the subsequent sensing/localization and communication designs.
II-A IPA Architecture
As shown in Fig. 1, each IPA can independently induce a certain phase shift (via the attached phase shifters in the polarformer) and amplitude change (via the attached attenuators in the polarformer) on the transmit/receive signals, thereby adjusting the antenna’s polarization. The phase shifters and attenuators in the polarformer enable precise polarization control by independently adjusting the phase and amplitude of signals. In this way, IPA can enable polarforming to exploit polarization diversity and adaptively control the antenna polarization for aligning polarization between transmitting and receiving antennas. We assume that each transmit/receive antenna consists of two orthogonally oriented linearly polarized elements, with one element for vertical polarization ($\mathcal{V}$-element) and the other for horizontal polarization ($\mathcal{H}$-element).
Specifically, each user $k$ is equipped with a single IPA, and the receive polarforming vector is denoted by
@@ II-A IPA Architecture
The BS has $B$ IPA subarrays, denoted by =$\mathcal{B}=\{1,2,\ldots,B\}$. Each IPA subarray consists of $N$ $(\geq 1)$ IPAs, denoted by $\mathcal{N}=\{1,2,\ldots,N\}$. The total number of transmit antennas at the BS is thus $BN$. Each IPA subarray is a uniform planar array (UPA) with a given size. All antennas within the same BS subarray share identical polarization characteristics determined by the propagation environment. Thus, their polarization can be controlled by the same polarforming vector. For the $b$-th IPA subarray at the BS, the transmit polarforming vector is given by
@@
where $\rho_{b,1}^{\mathrm{t}}\in[0,1]$ and $\rho_{b,2}^{\mathrm{t}}\in[0,1]$ represent the amplitude coefficients for the $\mathcal{V}$- and $\mathcal{H}$-elements, respectively, of each corresponding antenna at the $b$-th transmit IPA subarray. Similarly, $\psi_{b,1}^{\mathrm{t}}\in[0,2\pi)$ and $\psi_{b,2}^{\mathrm{t}}\in[0,2\pi)$ represent the phase shifts for the $\mathcal{V}$- and $\mathcal{H}$-elements, respectively, for the $b$-th transmit IPA subarray antennas. The phase and amplitude of the polarforming vector can be controlled either continuously or discretely. We consider the discrete control of amplitude and phase shifts for the ease of implementation. Let $Q_{\rho}$ and $Q_{\vartheta}$ denote the number of bits for polarforming amplitude and phase-shift control per antenna element, respectively. We thus have
Figure 3: Geometry illustration of the $b$-th IPA subarray at the BS and the IPA at user $k$.
@@
Each IPA can rotate and/or reposition within a given space. We assume that each user IPA can rotate while maintaining a fixed position. This is achieved with actuators or parasitic elements such as a rotating motor, or by manual adjustment, to enable physical movement. In contrast, all antennas on each BS IPA subarray can reposition and rotate together within a given convex 3D space $\mathcal{C}$ at the BS. This is mechanically controlled by connecting the subarrays to the central processing unit (CPU) of the BS via extendable and rotatable rods embedded with flexible wires, allowing the CPU to precisely control their 3D positions and rotations [9].
To facilitate the user/BS antenna movement description, we establish three Cartesian coordinate systems (CCSs). As shown in Fig. 3, the global CCS is denoted by $o\text{-}xyz$, with the BS’s CPU (i.e., BS center) centered at origin $o$. Each IPA subarray’s local CCS is denoted by $o^{\prime}\text{-}x^{\prime}y^{\prime}z^{\prime}$, with the subarray center as origin $o^{\prime}$. Each user’s local CCS is denoted by $o^{\prime\prime}\text{-}x^{\prime\prime}y^{\prime\prime}z^{\prime\prime}$, with origin $o^{\prime\prime}$ taken as the center of the user’s antenna. On the BS side, the position and rotation of the $b$-th IPA subarray, $b\in\mathcal{B}$, can be characterized by the following position vector and rotation vector
@@
where $x_{b}$, $y_{b}$ and $z_{b}$ represent the coordinates of the $b$-th IPA’s center in the global CCS; $\alpha_{b}$, $\beta_{b}$ and $\gamma_{b}$ all in $[0,2\pi)$ denote the rotation angles of the $b$-th IPA subarray with respect to (w.r.t.) the $x$-axis, $y$-axis, and $z$-axis in the global CCS, respectively. Given $\mathbf{u}_{b}$, the corresponding rotation matrix can be written as
@@
where $c_{x}=\cos(x)$ and $s_{x}=\sin(x)$ [9]. Let $\bar{\mathbf{r}}_{n}$ denote the position of the $n$-th antenna of the IPA subarray in its local CCS. Then, the position of the $n$-th antenna of the $b$-th IPA subarray in the global CCS can be expressed as

### v2

# Polarforming Antenna Enhanced Sensing and Communication: Modeling and Optimization

@@ Abstract
In this paper, we propose a novel polarforming antenna (PA) to achieve cost-effective wireless sensing and communication. Specifically, the PA can enable polarforming to adaptively control the antenna’s polarization electrically as well as tune its position/rotation mechanically, so as to effectively exploit polarization and spatial diversity to reconfigure wireless channels for improving sensing and communication performance. To analyze the performance gain of PA, we study an PA-enhanced integrated sensing and communication (ISAC) system that utilizes user location sensing to facilitate communication between an PA-equipped base station (BS) and PA-equipped users, by focusing on a new practical channel setup where the locations of users are nearly time-invariant but their orientations may change frequently (e.g., mobile phones rotated by spectators seated in a stadium while taking live photos). First, we model the PA channel in terms of transceiver antenna polarforming vectors and antenna positions/rotations. We then propose a two-timescale ISAC protocol, where in the slow timescale, user localization is first performed, followed by the optimization of the BS antennas’ positions and rotations based on the sensed user locations; subsequently, in the fast timescale, transceiver polarforming is adapted to cater to the instantaneous orientation of user devices in three-dimensional (3D) space, with the optimized BS antennas’ positions and rotations. We propose a new polarforming-based user localization method that uses a structured time-domain pattern of pilot-polarforming vectors to extract the common stable components in the PA channel across different polarizations based on the parallel factor (PARAFAC) tensor model. Moreover, we maximize the achievable average sum-rate of users by jointly optimizing the fast-timescale transceiver polarforming, including phase shifts and amplitude variations, along with the slow-timescale antenna rotations and positions at the BS. Simulation results validate the effectiveness of polarforming-based localization algorithm and demonstrate the performance advantages of polarforming, antenna placement, and their joint design in comparison with various benchmarks without polarforming or antenna position/rotation adaptation.
@@
Polarforming antenna (PA), polarforming optimization, antenna position/rotation optimization, integrated sensing and communication (ISAC), location sensing.
@@ I Introduction
To enable the transceiver’s full flexibility in antenna polarization and placement, we propose a novel polarforming antenna (PA), also called polarized six-dimensional movable antenna (6DMA) [9], as a new approach to improve the performance of wireless sensing and communication without the need to add more antennas and bear their additional cost and energy consumption. Specifically, an PA tunes the polarforming vector with adjustable signal amplitude and phase to control the polarization of each antenna at the transmitter and/or receiver, thus allowing dynamic adjustment of antenna polarization in response to real-time channel conditions and depolarization effects by exploiting polarization diversity. This is achieved through the electronically tunable polarformer, which contains phase shifters and attenuators that collectively adjust the phase and amplitude of the transmitted/received signal (see Fig. 1). Furthermore, each PA can be independently adjusted in terms of position and/or rotation to modify its polarization angle as well as adapt to the channel spatial distribution. This can be realized through mechanical control, which utilizes external mechanical structures with the aid of actuators, such as electric motors and precision gears, to achieve physical movement (see Fig. 1).
Figure 1: Illustration of the PA-enhanced sensing and communication system.
@@
It is worth noting that the PA system proposed in this paper differs significantly from our recently proposed 6DMA [9, 10, 11] and traditional polarization reconfigurable antenna [12]. Firstly, a mechanically controlled 6DMA, consisting of multiple three-dimensional (3D) rotatable and positionable antennas/subarrays, either lacks the antenna polarization control or is designed under a fixed polarization condition [13, 14, 15, 16, 17]. Secondly, although traditional polarization-reconfigurable antennas can adjust antenna polarization, they tune only the amplitude of the signal, which does not fully exploit polarization diversity. Moreover, polarization reconfigurable antennas with fixed antenna position/rotation cannot leverage antenna movement to achieve additional spatial DoFs. In contrast, employing the polarformer presented in [18], the proposed PA aims at collectively adjusting both the amplitude and phase of the signals to fully explore the benefits of polarization diversity, as well as adjusting its antenna position/rotation to enhance spatial DoFs. A summary of the above comparison is given in Table I.
@@
Besides enhancing communication performance, the proposed PA is to help improve wireless sensing/localization performance. Various wireless localization systems exist, such as those based on WLAN or cellular networks, which can determine the location of a user or target object using dedicated wireless reference signals between reference nodes and target nodes [1, 19, 20, 21]. A target object at an unknown location is often called agent node, while a reference node with a known position is referred to as anchor node. In some cases, the agent node determines its location by receiving reference signals from anchor nodes. However, there are other cases where the anchors receive reference signals from agents to localize the agents, and then inform them the estimated locations via a separate communication link [22]. Here, we consider the latter case, where multiple PAs are deployed at the anchor/base station (BS) to receive reference signals from single-PA users for user localization.
@@
Motivated by the above, we consider an PA-enhanced ISAC system that utilizes user location sensing to facilitate communication between PA-equipped users and an PA-equipped BS. In our considered system, the transmit and receive polarforming vectors, along with the positions and rotations of the PA subarrays at the BS, need to be jointly designed to achieve optimal sensing and communication performance. However, the design of these parameters depends on the user’s position and orientation. Thus, we consider a practical two-timescale channel setup, where the users’ positions change slowly and can be assumed to be fixed within a sufficiently long period termed as location coherence time, while their orientations may vary frequently during this period (intentionally or non-intentionally). This channel setup is applicable to many scenarios, such as spectators seated in a stadium watching a football match, where their smartphones remain stationary in position but are frequently rotated to take photos in different directions (see Fig. 2 (a)). Another example is for users playing virtual reality (VR) games at fixed locations, where their VR gadgets rotate frequently to enhance the gaming experience (see Fig. 2 (b)).
TABLE I: Comparison of PA with Other Related Technologies.
@@
Polarforming antenna (PA) High (Both phase and amplitude-adjustable polarforming) High (rotation and position) Electrical and mechanical Very high Medium
@@
In the proposed PA-enhanced ISAC system, the positions and rotations of antennas are mechanically controlled, which may have limited movement speeds, rendering them more suitable for adapting to slow-timescale channel variations, e.g., user locations. In contrast, transmit and receive polarforming of PA is electronically controlled, thus allowing their rapid adjustments to adapt to fast-timescale channel variations caused by, e.g., users’ frequent rotations. To design such a system, the acquisition of channel state information (CSI) is essential. However, since the BS can only measure user-BS channels at the positions and rotations where BS’s subarrays are physically deployed, it is difficult to obtain the instantaneous CSI across the entire antenna movement region at the BS and for arbitrary users’ positions and rotations. Fortunately, in our considered scenarios (see Fig. 2), the user’s locations remain unchanged within each location coherence time, which can be acquired and used for the optimization of BS antenna positions/rotations. Then, with the optimized antenna positions/rotations at the BS, the instantaneous CSI of users due to their rotations can be measured for designing the fast-timescale transmit and receive polarforming. The main contributions of this paper are summarized as follows.
@@
  • First, we propose the new PA to electrically adjust antenna polarization through polarforming vectors to align the transmitter and receiver polarizations, as well as mechanically adjust antenna position/rotation to adapt to user spatial channel distribution. Moreover, we model the PA channel and decouple the channel into stable unpolarformed components and dynamic polarformed components, which provides essential insights into our considered channel setup and facilitates the determination of parameters in the PA-enhanced system.
@@
  • Next, we study an PA-enhanced ISAC system that utilizes user location sensing to facilitate communications between PA-equipped users and an PA-equipped BS. For the new practical channel setup where user locations are nearly static but orientations may vary frequently, we propose a two-timescale ISAC protocol. In the slow timescale, user localization is first performed, based on which the BS antennas’ positions and rotations are optimized, while subsequently in the fast timescale, transceiver polarforming is adapted to cater to the instantaneous user/device rotations in 3D space.
@@
  • Moreover, we design a new polarforming-based user localization method that leverages controllable polarforming vectors to create a time-domain pattern of pilot-polarforming vectors. This enables the extraction of stable line-of-sight (LoS) components in the PA channel across different polarizations using the parallel factor (PARAFAC) tensor. Thereby, user locations can be accurately determined from the estimated LoS components. Simulation results validate the effectiveness of the proposed localization method by exploiting the new DoF of polarforming.
@@
  • Finally, we formulate an optimization problem to maximize the weighted sum-rate of users by jointly designing transceiver polarforming and BS antenna positions/rotations based on sensed user locations. To solve this non-convex problem, we propose an efficient algorithm that accounts for PA movement constraints, discrete polarforming amplitude and phase shifts, and enables efficient variable updates via closed-form solutions or simple iterations. We show by simulation that polarforming optimization improves the achievable rate over fixed polarization, with more gains achieved when being jointly optimized with antenna position/rotation.
@@
Figure 2: Examples of user rotation scenarios in PA-enhanced ISAC systems.
@@
The rest of this paper is organized as follows. Section II presents the PA architecture and the corresponding channel model. Section III presents a two-timescale protocol design for the PA-enhanced ISAC system. Section IV presents the polarforming-based localization algorithm. Section V addresses antenna polarforming and position/rotation optimization. Section VI presents simulation results for performance evaluation. Finally, Section VII concludes this study.
@@ II System Model
We consider an PA-enhanced ISAC system as shown in Fig. 1, where multiple PAs are deployed at the BS to receive reference signals from single-PA users for user localization as well as to communicate with these users over a given frequency band. Without loss of generality, we consider the downlink communications from the BS to users. In this section, we first describe the new PA architecture and then present its corresponding channel model, which lays the foundation for the subsequent sensing/localization and communication designs.
II-A PA Architecture
As shown in Fig. 1, each PA can independently induce a certain phase shift (via the attached phase shifters in the polarformer) and amplitude change (via the attached attenuators in the polarformer) on the transmit/receive signals, thereby adjusting the antenna’s polarization. The phase shifters and attenuators in the polarformer enable precise polarization control by independently adjusting the phase and amplitude of signals. In this way, PA can enable polarforming to exploit polarization diversity and adaptively control the antenna polarization for aligning polarization between transmitting and receiving antennas. We assume that each transmit/receive antenna consists of two orthogonally oriented linearly polarized elements, with one element for vertical polarization ($\mathcal{V}$-element) and the other for horizontal polarization ($\mathcal{H}$-element).
Specifically, each user $k$ is equipped with a single PA, and the receive polarforming vector is denoted by
@@ II-A PA Architecture
The BS has $B$ PA subarrays, denoted by =$\mathcal{B}=\{1,2,\ldots,B\}$. Each PA subarray consists of $N$ $(\geq 1)$ PAs, denoted by $\mathcal{N}=\{1,2,\ldots,N\}$. The total number of transmit antennas at the BS is thus $BN$. Each PA subarray is a uniform planar array (UPA) with a given size. All antennas within the same BS subarray share identical polarization characteristics determined by the propagation environment. Thus, their polarization can be controlled by the same polarforming vector. For the $b$-th PA subarray at the BS, the transmit polarforming vector is given by
@@
where $\rho_{b,1}^{\mathrm{t}}\in[0,1]$ and $\rho_{b,2}^{\mathrm{t}}\in[0,1]$ represent the amplitude coefficients for the $\mathcal{V}$- and $\mathcal{H}$-elements, respectively, of each corresponding antenna at the $b$-th transmit PA subarray. Similarly, $\psi_{b,1}^{\mathrm{t}}\in[0,2\pi)$ and $\psi_{b,2}^{\mathrm{t}}\in[0,2\pi)$ represent the phase shifts for the $\mathcal{V}$- and $\mathcal{H}$-elements, respectively, for the $b$-th transmit PA subarray antennas. The phase and amplitude of the polarforming vector can be controlled either continuously or discretely. We consider the discrete control of amplitude and phase shifts for the ease of implementation. Let $Q_{\rho}$ and $Q_{\vartheta}$ denote the number of bits for polarforming amplitude and phase-shift control per antenna element, respectively. We thus have
Figure 3: Geometry illustration of the $b$-th PA subarray at the BS and the PA at user $k$.
@@
Each PA can rotate and/or reposition within a given space. We assume that each user PA can rotate while maintaining a fixed position. This is achieved with actuators or parasitic elements such as a rotating motor, or by manual adjustment, to enable physical movement. In contrast, all antennas on each BS PA subarray can reposition and rotate together within a given convex 3D space $\mathcal{C}$ at the BS. This is mechanically controlled by connecting the subarrays to the central processing unit (CPU) of the BS via extendable and rotatable rods embedded with flexible wires, allowing the CPU to precisely control their 3D positions and rotations [9].
To facilitate the user/BS antenna movement description, we establish three Cartesian coordinate systems (CCSs). As shown in Fig. 3, the global CCS is denoted by $o\text{-}xyz$, with the BS’s CPU (i.e., BS center) centered at origin $o$. Each PA subarray’s local CCS is denoted by $o^{\prime}\text{-}x^{\prime}y^{\prime}z^{\prime}$, with the subarray center as origin $o^{\prime}$. Each user’s local CCS is denoted by $o^{\prime\prime}\text{-}x^{\prime\prime}y^{\prime\prime}z^{\prime\prime}$, with origin $o^{\prime\prime}$ taken as the center of the user’s antenna. On the BS side, the position and rotation of the $b$-th PA subarray, $b\in\mathcal{B}$, can be characterized by the following position vector and rotation vector
@@
where $x_{b}$, $y_{b}$ and $z_{b}$ represent the coordinates of the $b$-th PA’s center in the global CCS; $\alpha_{b}$, $\beta_{b}$ and $\gamma_{b}$ all in $[0,2\pi)$ denote the rotation angles of the $b$-th PA subarray with respect to (w.r.t.) the $x$-axis, $y$-axis, and $z$-axis in the global CCS, respectively. Given $\mathbf{u}_{b}$, the corresponding rotation matrix can be written as
@@
where $c_{x}=\cos(x)$ and $s_{x}=\sin(x)$ [9]. Let $\bar{\mathbf{r}}_{n}$ denote the position of the $n$-th antenna of the PA subarray in its local CCS. Then, the position of the $n$-th antenna of the $b$-th PA subarray in the global CCS can be expressed as